OpenStudy (anonymous):

Let n be an odd integer with 11 positive divisors. find the number of positive divisors of 8n^3.

ganeshie8 (ganeshie8):

Hey!

OpenStudy (anonymous):

Hello!

OpenStudy (amistre64):

how do we describe 11 positive divisors? say we have a number like 30, what are its positive divisors?

OpenStudy (amistre64):

are they: 2,3,5 ? or, 1,30, 2,15, 3,10, 5,6 ??

OpenStudy (anonymous):

the second set of numbers you gave would be it's positive divisors.

OpenStudy (anonymous):

so number of divisors is tau function right ? are you familiar with it ?

OpenStudy (anonymous):

so when you say n is an odd integer with 11 positive divisors means \(\Large \tau(n)=11\)

OpenStudy (anonymous):

now tau function is multiplicative which means it have this property \(\Large \tau(a\times b)=\tau(a)\times\tau(b)\)

OpenStudy (anonymous):

hello there @contradiction are u with me ?

OpenStudy (anonymous):

so \(\Large \tau(8n^3)=\tau(8)\times(\tau(n))^3\)

OpenStudy (anonymous):

oh!

OpenStudy (anonymous):

sorry for the late response

OpenStudy (amistre64):

dont divisors some in groups of 2?

OpenStudy (amistre64):

hmm, i spose if one set was a perfect square then the grouping a,a would represent a single divisor as opposed to 2 of them

OpenStudy (amistre64):

an algorithm online says that the product of the exponents of the prime factorization ... when you add 1 gives us the number of divisors. so like: n=3^(10) has 11 divisors: (10+1)(0+1)

OpenStudy (amistre64):

n^3 = 3^(30) 2^3 = 8, so (3+1)(30+1) seems to be a specific approach

OpenStudy (anonymous):

ah, right, that makes sense

OpenStudy (anonymous):

okay, i got it! thank you!